# 2Università degli Studi di Siena, Via Roma 56, Siena, Italy

### v1

@@ 6 Conclusions
In recent years, the improving sensitivity of -ray telescopes has led to the identification of several YMSCs as -ray sources. Nonetheless, nowadays, the detection of individual YMSCs still remains a challenging task due to their low surface brightness. This implies that most of the -ray flux from YMSCs is likely detected as an unresolved contribution to the diffuse Galactic -ray emission. Indeed, recent analyses of the diffuse -ray emission have shown a prominent excess, likely related to unresolved sources, ranging from a few tens of GeV up to hundreds of TeV. In this work, we estimated the contribution of YMSCs to the diffuse emission by simulating the entire Galactic populations of YMSCs, starting from observed properties of local clusters. We modeled the -ray spectrum of each synthetic YMSC considering a purely hadronic scenario and assuming that particle acceleration occurs at the collective cluster wind termination shock. We found that the contribution of the cluster population to the diffuse emission is sizeable above few hundreds of GeV in the inner region of the Milky Way when the particle transport in the wind-blown bubble is governed by Kolmogorov or Kraichnan type diffusion. In the outer region of the Milky Way we could not draw any solid conclusion due to the large statistical uncertainties related to the realisation of the YMSC population. A more reliable estimate in this region should be made by directly using existing YMSC surveys.
@@ Acknowledgements.
- Blasi & Morlino (2023) Blasi, P. & Morlino, G. 2023, arXiv e-prints, arXiv:2307.11663
@@
- Menchiari et al. (2024) Menchiari, S., Morlino, G., Amato, E., Bucciantini, N., & Beltrán, M. T. 2024, submitted for publication to A&A
@@ Appendix A Generation of the mock stellar population
Appendix B Radial distribution of YMSCs
1.
2.
Parameters | Arm 1 | Arm 2 | Arm 3 | Arm 4 | Local Spur⋆ |
Ri [kpc] | 3.27 | 4.29 | 3.58 | 3.98 | 8.16 |
[∘] | 9.87 | 10.51 | 10.01 | 8.14 | 2.71 |
[∘] | 38.5 | 189 | 215.2 | 320.1 | 50.6 |
On top of the spiral structure, we also take into account the structure of the innermost region. Here we consider the presence of the Galactic bar and the Near 3 kpc and Far 3 kpc arms. The first is modeled as an ellipse with semi-major axis of 3.3 kpc and an aspect ratio of 4:10 (Churchwell et al. 2009). The Near/Far-3-kpc Arms are modeled using an elliptical annulus, with a semi-major axis of 4.1 kpc, and an aspect ratio of 0.54 (a semi-minor axis of 2.2 kpc) (Green et al. 2011).
Spiral Arm to 4, if ;
Local spur if and ;
Spiral Arm or Near/Far-3-kpc Arms and Galactic Bar, if kpc;
Near/Far-3-kpc Arms and Galactic Bar, if kpc.
Once all clusters are associated to a structure, we proceed to perturb their positions according to a Gaussian distribution, following the fundamental concept that both Spiral Arms and the Near and Far structure possess an intrinsic thickness. The width of the spiral arm increases with galactocentric distance, so that the probability of finding a source displaced from the center of the arm by , is (Faucher-Giguère & Kaspi 2006):
As for the clusters belonging to the Near and Far arms, we extract from a Gaussian probability distribution the Galactic x and y coordinates with a spread of kpc, such that the probability at returns a scattering compatible with the observed radial thickness of 0.5 kpc (Green et al. 2011). Notice that the thickness of a spiral arm in YMSCs may be slightly different from that inferred for pulsars, however the result on the diffuse emission depends very weakly on this parameter.
Finally, we generate the vertical coordinate () following the observed gas distribution, i.e. an exponential profile with a characteristic spread of 100 pc (Strong et al. 2000):
Appendix C Definition of masks
@@
We used three different kinds of masks. The first mask, M, is designed to remove regions of the Galactic Plane where, on average, most of the galactic sources are located. Following the shape of M, one can easily see that, for , a large fraction of the Galactic Plane is masked. We hence define M so as to remove the zone of the sky defined between and . We also remove the region corresponding to the projected size of the Local Spiral arm, as an over-density of sources is expected there. Such a region corresponds to a circular area centered on with radius . Notice that in ROI2 a sparse distribution of the galactic sources is expected, as the outer region of the Milky Way is less crowded. Consequently, there is not a specific zone of the plane to mask. Hence, we define this specific type of mask only for ROI1.
The second type of masks, M and M, is defined so as to remove the emission from those YMSCs that would be detectable by the LHAASO experiment. To do so, we mask the positions of the sky where the significance of the surface brightness at 100 TeV () exceeds by 5 the brightness of the measured galactic diffuse emission, considered here as the nominal background. We use the emission at 100 TeV as it correspond to the energy where LHAASO has the best sensitivity (Cao et al. 2019). Hereafter, the subscript ”100” indicates that the quantity is evaluated at 100 TeV. At each position of the sky belonging to the ROIs with coordinates (, ), the significance of the emission is calculated as:
@@
In Equation (C) the angles and are the equatorial declination of the LHAASO site and the given position of the sky, respectively. Finally, the last type of mask, M, is obtained by merging M with M. In ROI1 the M mask is the one that best reproduces M.
Figure 4 shows an example of how M, M and M compare with M. This specific configuration refers to the Galaxy realization showed in Figure 3 with the assumption of Kraichnan diffusion and inclusion of the WR contribution. Using all the 100 different realisations of the Milky Way, we can quantitatively estimate the difference between our masks and the original ones by evaluating the median ratio of the non-masked areas. For the M mask we found this ratio to be 1.22, 1.19 and 1.1 for the Kolmogorov, Kraichnan and Bohm cases, respectively. The same ratios calculated for the ROI2 are 1.24, 1.24 and 1.22. Those numbers are estimated including WR stars in the stellar population. We found no significant difference when WRs are excluded. In summary, our sky regions is between 10% and 20% larger than the one considered by Cao et al. (2023a).
The same method employed to generate the masks M and M can also be used to estimate the number of YMSCs that should have been detected by LHAASO. This can be seen in Figure 5 were the distribution of the number of detected clusters over the 100 different realisations is shown. Note that this number of detections must be considered as an order of magnitude estimate, as the detectability of a cluster should be evaluated after considering the instrument sensitivity over the entire energy range. Interestingly, the number of detections for the Kolmogorov and Kraichnan cases is consistent with the number of sources in the LHAASO catalog that can be associated to SCs (Mitchell et al. 2024). In contrast, the predicted median number of YMSCs in the Bohm case ( 90 if WRs are included in the stellar populations) is of the same order of the total number of sources in the LHAASO catalog, suggesting that this scenario is likely to be too extreme.
Appendix D Diffuse emission from the CR sea

### v2

@@ 6 Conclusions
In recent years, the improving sensitivity of -ray telescopes has led to the identification of several YMSCs as -ray sources. Nonetheless, nowadays, the detection of individual YMSCs still remains a challenging task due to their low surface brightness. This implies that most of the -ray flux from YMSCs is likely detected as an unresolved contribution to the diffuse Galactic -ray emission. Indeed, recent analyses of the diffuse -ray emission have shown a prominent excess, likely related to unresolved sources, ranging from a few tens of GeV up to hundreds of TeV. In this work, we estimated a lower limit to the contribution of YMSCs to the diffuse emission by simulating the entire Galactic populations of YMSCs, starting from observed properties of local clusters. We modeled the -ray spectrum of each synthetic YMSC considering a purely hadronic scenario and assuming that particle acceleration occurs at the collective cluster wind termination shock. We found that the contribution of the cluster population to the diffuse emission is sizeable above few tens of GeV in the inner region of the Milky Way when the particle transport in the wind-blown bubble is governed by Kolmogorov or Kraichnan type diffusion. In the outer region of the Milky Way we could not draw any solid conclusion due to the large statistical uncertainties related to the realisation of the YMSC population. A more reliable estimate in this region should be made by directly using existing YMSC surveys.
@@ Acknowledgements.
- Blasi & Morlino (2024) Blasi, P. & Morlino, G. 2024, MNRAS, 533, 561
@@
- Lada & Lada (2003) Lada, C. J. & Lada, E. A. 2003, ARA&A, 41, 57
- Lancaster et al. (2021) Lancaster, L., Ostriker, E. C., Kim, J.-G., & Kim, C.-G. 2021, ApJ, 914, 90
@@
- Menchiari et al. (2024) Menchiari, S., Morlino, G., Amato, E., Bucciantini, N., & Beltrán, M. T. 2024, A&A, 686, A242
@@
- Niedzielski & Skorzynski (2002) Niedzielski, A. & Skorzynski, W. 2002, Acta Astron., 52, 81
@@
- Silich & Tenorio-Tagle (2013) Silich, S. & Tenorio-Tagle, G. 2013, ApJ, 765, 43
@@ Appendix A Generation of the mock stellar population
Appendix B Definition of masks
@@
We used three different kinds of masks. The first mask, , is designed to remove regions of the Galactic Plane where, on average, most of the galactic sources are located. Following the shape of M, one can easily see that, for , a large fraction of the Galactic Plane is masked. We hence define so as to remove the zone of the sky defined between and . We also remove the region corresponding to the projected size of the Local Spiral arm, as an over-density of sources is expected there. Such a region corresponds to a circular area centered on with radius . Notice that in ROI2 a sparse distribution of the galactic sources is expected, as the outer region of the Milky Way is less crowded. Consequently, there is not a specific zone of the plane to mask. Hence, we define this specific type of mask only for ROI1.
The second type of masks, , is defined so as to remove the emission from those YMSCs that would be detectable by the LHAASO experiment. To do so, we mask the positions of the sky where the significance of the surface brightness at 100 TeV () exceeds by 5 the brightness of the measured galactic diffuse emission, considered here as the nominal background. We use the emission at 100 TeV as it correspond to the energy where LHAASO has the best sensitivity (Cao et al. 2019). Hereafter, the subscript ”100” indicates that the quantity is evaluated at 100 TeV. At each position of the sky belonging to the ROIs with coordinates (, ), the significance of the emission is calculated as:
@@
In Equation (B) the angles and are the equatorial declination of the LHAASO site and the given position of the sky, respectively. Finally, the last type of mask, , is obtained by merging with . In ROI1 the mask is the one that best reproduces M. Figure 4 shows how the contribution to the diffuse gamma-ray flux in ROI1 changes after applying the (left column), (central column) and (right column) masks.
Figure 5 shows an example of how , , and compare with M. This specific configuration refers to the Galaxy realization showed in Figure 2 with the assumption of Kraichnan diffusion and inclusion of the WR contribution. Using all the 100 different realisations of the Milky Way, we can quantitatively estimate the difference between our masks and the original ones by evaluating the median ratio of the non-masked areas. For the mask we found this ratio to be 1.22, 1.19 and 1.1 for the Kolmogorov, Kraichnan and Bohm cases, respectively. The same ratios calculated for the ROI2 are 1.24, 1.24 and 1.22. Those numbers are estimated including WR stars in the stellar population. We found no significant difference when WRs are excluded. In summary, our sky regions is between 10% and 20% larger than the one considered by Cao et al. (2023a).
The same method employed to generate the mask can also be used to estimate the number of YMSCs that should have been detected by LHAASO. This can be seen in Figure 6 were the distribution of the number of detected clusters over the 100 different realisations is shown. Note that this number of detections must be considered as an order of magnitude estimate, as the detectability of a cluster should be evaluated after considering the instrument sensitivity over the entire energy range. Interestingly, the number of detections for the Kolmogorov and Kraichnan cases is consistent with the number of sources in the LHAASO catalog that can be associated to SCs (Mitchell et al. 2024). In contrast, the predicted median number of YMSCs in the Bohm case ( 100 if WRs are included in the stellar populations) is of the same order of the total number of sources in the LHAASO catalog, suggesting that this scenario is likely to be too extreme.
Appendix C Diffuse emission from the CR sea